Suppose a portfolio is delta-neutral and has a gamma of -5600. A call option trades with a delta of 0.5 and a gamma of 0.8. You buy this option to gamma-hedge. How many shares of stock do you need to buy (sell) to delta-hedge your portfolio after you added the option?

a. Sell 3500 shares

b. Buy 5000 shares

c. Buy 3500 shares

d. Sell 5000 shares
Answer #1

Delta of Existing POrtfolio = 0 (Neutral)

Gamma of Existing POrtfolio = - 5600

Delta of Call option = 0.50

Gamma of Call option = 0.80

Since we have 100 shares in one contract

Delta Position of Call Option = 0.50 * 100 = 50

Gamma Position of Call Option = 0.80 * 100 = 80

Call Option we need to buy to gamma hedge the portfolio = 5600 / 80 = 70

Therefore to to delta hedge the new portolio we need to sell = 70 * 50 * 1 = 3500 shares

Option A is correct.